OpenStudy (anonymous):

Which of the following is the ratio of circumference to diameter? radius center circle pi

OpenStudy (solomonzelman):

Circumference = 2πr Diameter = 2π Circumference : Diameter → 2πr : 2π → r:1 .

OpenStudy (solomonzelman):

So, which one do you think it is ?

OpenStudy (anonymous):

..... im sorry idk math is like a different language to me can u explain it in a different way maybe?

OpenStudy (anonymous):

a or D?

OpenStudy (solomonzelman):

Okay, lets see.... Do you understand what I mean by saying `Circumference = 2πr` and `Diameter = 2π` ?

OpenStudy (anonymous):

not really

OpenStudy (solomonzelman):

How about `Circumference = 2 × π × radius` and `Diameter = 2 × π `

OpenStudy (anonymous):

ok but how do i do that?

OpenStudy (solomonzelman):

So, the circumference = 2 × π × radius Diameter = 2 × π So the ratio of circumference to diameter (in a fraction) would be Circumference / Diameter and this is convert-able to 2 × π × radius ------------ 2 × π `2 × π` cancels, and you get that the ratio of circumference to diameter is equivalent to the radius.

OpenStudy (anonymous):

OoO ok....

OpenStudy (anonymous):

I still have no clue what it is, a?

OpenStudy (solomonzelman):

Yes !

OpenStudy (solomonzelman):

I don't know what I just clicked, but I copy pasted the previous reply

OpenStudy (anonymous):

ha wow thank couldn't have done it without you!!!

OpenStudy (solomonzelman):

Anytime:) Have a good one !